solve the proportion
5/ x-1= 7/x

Answer by checkley77(12844)

5/ x-1= 7/x cross multiply
7(x-1)=5x
7x-7=5x
7x-5x=7
2x=7
x=7/2
x=3.5 ans.
Proof:
5/(3.5-1)=7/3.5
5/2.5=2
2=2